OpenStudy (anonymous):

find x intercepts if any for the graph of the quadratic function f(x)=(x+2)^2-4

OpenStudy (zzr0ck3r):

\((x+2)^2-4=0\) can you solve that?

OpenStudy (anonymous):

x=0

OpenStudy (zzr0ck3r):

and?

OpenStudy (anonymous):

when you plug in your zero it is it is 0=0

OpenStudy (zzr0ck3r):

there is more than one

OpenStudy (zzr0ck3r):

but you need to solve it

OpenStudy (zzr0ck3r):

\((x+2)^2-4=0\implies (x+2)^2=4\implies x+2=\pm\sqrt{4}\implies x=\pm 2-2\) So \(x=-2-2\) and \(x=2-2\)

OpenStudy (anonymous):

yes I see thanks